OpenStudy (anonymous):

a + 3c − 4b − c + a − 2b

mathslover (mathslover):

first of all combine like terms:

mathslover (mathslover):

\[\large{\color{blue}{a+3c-4b-c+a-2b}}\] \[\large{\color{red}{a+a+3c-c-4b-2b}}\]

mathslover (mathslover):

first of all solve for a + a can you tell me what is a + a ?

mathslover (mathslover):

@davisgurl101 . please reply

OpenStudy (anonymous):

2s

OpenStudy (anonymous):

2a

mathslover (mathslover):

good

OpenStudy (anonymous):

= a + 3c − 4b − c + a − 2b = (a + a) + (-4b - 2b) + (3c - c) = 2a - 6b + 2b

OpenStudy (anonymous):

2a - 6b + 2c do you know how to get this?

mathslover (mathslover):

now we have : \[\large{2a+3c-c-4b-2b}\] what is 3c-c ? @davisgurl101

OpenStudy (anonymous):

sure